OpenStudy (anonymous):

PLEASE HELP: How many gallons of a 30% acid solution must be mixed with 50 gallons of a 70% acid solution to make 55% acid solution? Explain.

OpenStudy (anonymous):

Let x be gallons of 30% acid solution. Total gallons=x+50 0.3x+0.7(50)=(x+50)(0.55) Solve for x.